Determine which shipping alternative would be most economical to ship 69 boxes of parts when each box has a price of $298 and holding costs are 28 percent of price, given this shipping information: overnight, $435, two-day, $295, six-day, $185. (Round your intermediate calculations to 2 decimal places.)

  • Overnight

  • Six-day

  • Two-day
Answer #1

Holding cost for 69 boxes = 28% of 69*298 = 83.44 = 5757.36

Per day holding cost for 69 boxes = 5757.36/365 = 15.77

Total cost for Overnight shipping =435+15.77*1 = 450.77


Total cost for 2-day shipping= 295+2*15.77 = 326.54


Total cost for 6-day shipping= 185+6*15.77 = 279.62

Correct answer is Six-day as Six-day shipping incurs least cost